OpenStudy (unimatix):

Compute the standard parts: (sqrt(H+1))/(sqrt(2H)+sqrt(H-1)), H positive infinite

OpenStudy (solomonzelman):

\(\large\color{blue}{\displaystyle\lim_{x \rightarrow ~\infty}\frac{\sqrt{x+1}}{\sqrt{2x}+\sqrt{x-1}}}\) can I interpret your question like this?

OpenStudy (unimatix):

I believe that would be the same. Kind of trying to teach myself. The answer is given as 1 / ( 1 + sqrt(2) ) OR sqrt(2) - 1. I don't know how to solve it though.

OpenStudy (solomonzelman):

rationalize the denominator

OpenStudy (welshfella):

yes

OpenStudy (triciaal):